The perimeter of a parallelogram is 60. If one side is 5 longer than the side next to it, what are the lengths of the sides?
Answer by 2897696(96)

length(longest side): x+5
width: x
perimeter of parallelogram is 60
x+5+x+5+x+x=60
4x+10=60
4x=50
x=12.5<-width
length: x+5 -> 12.5 + 5=17.5

side length: width 12.5 &
length: 17.5